# Ab initio recombination in the expanding ultracold plasmas

### v1

@@ Abstract
The efficiency of recombination is of crucial importance for the existence of ultracold plasmas, particularly, the ones formed in the magneto–optical traps. Unfortunately, a straightforward simulation of the recombination encounters the problem of huge difference in the spatial and temporal scales for free and bound motion of the electrons. As a result, only the “virtual” electron–ion pairs are usually reproduced in such simulations, and it is necessary to employ some additional criteria to identify them with the recombined atoms (this might be a minimal number of revolutions of the electron about the nearest ion or a maximal distance between them). It is the aim of this paper to present the first successful ab initio simulation of the recombination without any auxiliary assumptions. We employed a special algorithm, which was based on: (i) using the “scalable” reference frame, co-moving with the expanding plasma, (ii) dynamical choice of the number of “mirror” cells, taking into account in calculation of the Coulomb sums, and (iii) accurate treatment of the singular interparticle interactions, without any truncation or “softening” of the Coulomb forces. Then, the recombination events are identified by a series of sharp equidistant peaks in the kinetic and/or potential energies for a sample of particles, which are caused by the captured electrons passing near the pericenters of their orbits; and this is confirmed by a detailed inspection of the particle trajectories. Thereby, we were able to trace formation of the real—rather than “virtual”—electron–ion pairs. The total efficiency of recombination for the realistic experimental conditions was found to be about 20%, which is in perfect agreement both with the laboratory measurements and with the earlier semi-empirical simulations.
@@ I Introduction
The ultracold plasmas (UCP), formed after release of the bunches of photoionized atoms from the magneto-optical traps Killian et al. (1999), represent a relatively new branch of plasma physics, emerged in the very late 1990s and early 2000s Gould and Eyler (2001); Bergeson and Killian (2003); Killian (2007); Killian et al. (2007). The possibility of existence of such plasmas for an appreciable time intervals crucially depends on the efficiency of recombination, since it might be intuitively expected that the charged particles with a sufficiently strong Coulomb coupling will quickly collapse into the neutral pairs.
@@
It is interesting to mention that the corresponding numerical simulations started already in the early 1990’s, i.e., a decade before the first experiments with UCP. Surprisingly, these simulations demonstrated that the ultracold plasmas experienced a very small recombination, i.e., could survive for a sufficiently long time Mayorov, Tkachev, and Yakovlenko (1994). Unfortunately, it was suspected by some opponents that the anomalous resistance of charged particles with respect to the recombination might be just an artifact of the particular numerical methods. For example, the recombined neural pairs could be efficiently destroyed by the collisions with walls of the computational box Ignatov et al. (1995); Mayorov, Tkachev, and Yakovlenko (1995), if the reflective boundary conditions are imposed there. Moreover, the anomalously small recombination rate might be caused just by the accumulation of numerical errors, preventing formation of the bound electron–ion pairs. This problem is especially severe because of the huge difference in the characteristic temporal and spatial scales for the free and bound motion of the charged particles in plasmas.
@@
In view of the above-mentioned problems, the most of works on recombination in UCP employed a semi-empirical two-step procedure: Firstly, the electron distribution function was simulated numerically by the molecular-dynamic approach, and then it was used for analytical estimates of the recombination rate, e.g., employing the idea of electron diffusion in the energetic space Bobrov et al. (2011).
@@
On the other hand, as regards the ab initio modeling of recombination, the most of works employed some artificial criteria for the formation of bound electron–ion pairs: For example, it was assumed that the act of recombination took place if an electron performed a certain number of revolutions about the nearest ion (e.g., 4 complete revolutions) Lankin and Norman (2009a). Yet another approach was to assume that all electrons located closer than a specified distance from the nearest ions (e.g., 20% of the Wigner–Seitz radius) will ultimately recombine Niffenegger, Gilmore, and Robicheaux (2011). Such artificial criteria were necessary because the commonly-used numerical schemes did not allow ones to get the sufficiently stable electron–ion pairs: they were destroyed rather quickly due to accumulation of the computational errors.
@@
It is the aim of the present work to describe a specially designed numerical algorithm that is really able to trace formation of the stable electron–ion pairs and, therefore, to serve as a testbed for other methods. In fact, this algorithm closely resembles the one employed in our earlier paper Dumin (2011), which was aimed at ab initio simulation of the electron temperature evolution in ultracold plasma bunches experiencing a considerable expansion. The simulated dependence $T_{e}\propto t^{\alpha}$ with exponent $\alpha\in[-1.25,-1.08]$ turned out to be in perfect agreement with the results of experiment Fletcher, Zhang, and Rolston (2007), where $\alpha=-(1.2{\pm}0.1)$. It is important to emphasize that this coincidence was achieved exactly from “the first principles”, without postulation of any additional heat sources or sinks. As will be seen below, the same numerical scheme provides also very good results in the simulation of recombination.
@@ II.1 Geometrical configuration
Figure 1: Sketch of calculation of the Coulomb sums, where each electron $e$ interacts not only with a specified ion $i$ in the basic cell (solid arrow) but also with infinite number of its mirror images $i^{\prime}$ (dashed arrows).
We shall consider an infinite volume of plasma expanding uniformly in all directions. It is assumed to be composed of the identical cubic cells filled with some number of ions and the same number of electrons, as illustrated in Fig. 1. Then, the Coulomb forces are calculated between all particles in this system i.e., not only within the same basic cell but also with all its “mirror” images:
@@
$+$ $\sum_{k=1}^{3}\>\sum_{n_{k}=-\infty,\,n_{k}\neq 0}^{+\infty}\Bigg\{\sum_{j=1}^{N}e^{2}\frac{[\,{\bf R}_{j}+L\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]-{\bf r}_{i}}{\big|[\,{\bf R}_{j}+L\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]-{\bf r}_{i}\big|^{3}}$
$+$ $\sum_{j=1}^{N}\,e^{2}\frac{{\bf r}_{i}-[\,{\bf r}_{j}+L\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]}{\big|{\bf r}_{i}-[\,{\bf r}_{j}+L\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]\big|^{3}}\Bigg\}\>.$
Here, ${\bf F}_{i}$ if the force acting on the $i$’th electron, ${\bf r}_{j}$ and ${\bf R}_{j}\;(j=1,\dots,N)$ are the electronic and ionic coordinates, $N$ is the number of charged particles of each kind in the basic cell, $e$ is the electron charge, $L$ is the linear size of the basic cell, ${\bf e}_{l}\;(l=1,2,3)$ are the unitary vectors of the Cartesian coordinate system, and subscripts $n_{k}\;(k=1,2,3)$ specify the number of the mirror cell in each direction. We do not write here a similar formula for the forces acting on ions: since the ions are much more massive than electrons, they can be usually assumed to move by inertia (some additional explanations will be given below).
@@
In fact, summation in formula (1) is performed over a finite number of mirror cells until the specified accuracy of convergence of the Coulomb forces is achieved. These mirror cells are arranged into the cubic shells of increasing radii around the basic cell; see Fig. 2. Therefore, although we used a relatively small number of charged particles in the basic cell (e.g. a few dozens or a few hundreds), the total number of particles taken into account in calculation of the Coulomb sums becomes huge. For example, if the relative accuracy of calculation is specified to be about $10^{-5}$, then even at $N=10$ the total number of particles involved in the computation turns out to be from a few hundred thousand to a few million. Let us emphasize that this number is not fixed but chosen dynamically at each step of integration.
Of course, a possible alternative to the straightforward summation over the cells, might be the well-known Ewald procedure, whose original idea can be found, e.g., in monograph Ziman (1972) and more elaborated versions in paper Demyanov and Levashov (2022). However, since our original intention was to follow the ab initio approach as strictly as possible, we did not use here the Ewald approximation.
II.2 The scalable reference frame
Yet another serious problem in simulation of the expanding plasma clouds is a considerable change in the spatial scale of the system—and, therefore, in the Coulomb forces—in the course of time, which makes it difficult to keep the accuracy of integration at a stable level. To get around this obstacle, we introduce the scalable reference frame, expanding with the average expansion rate of the plasma. Namely, a length of the basic cell is assumed to increase as
$L(t)=L_{0}+u_{0}\,t\>,$ (2)
@@ II.2 The scalable reference frame
In other words, the simulated particles perform their motion in a box of the fixed size (7), thereby mitigating the problem of very large variation of the spatial coordinates in the case of considerable plasma expansion. As usual, the periodic boundary conditions are imposed at the box boundaries: when one of the particles leaves the box through one of its faces, the identical particle with the same velocity enters the box through the opposite face.
The classical equation of motion of the $i$’th electron,
$m_{e}\,d^{2}{\rm\bf r}_{i}/dt^{2}=\,{\rm\bf F}_{i}\,$, (8)
in the scale-dependent dimensionless variables (4) and (5) is reduced to
@@
$+$ $\sum_{k=1}^{3}\>\sum_{n_{k}=-\infty,\,n_{k}\neq 0}^{+\infty}\Bigg\{\sum_{j=1}^{N}\frac{[\,{\bf R}^{*}_{j}+L^{*}\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]-{\bf r}^{*}_{i}}{\big|[\,{\bf R}^{*}_{j}+L^{*}\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]-{\bf r}^{*}_{i}\big|^{3}}$
$+$ $\sum_{j=1}^{N}\,\frac{{\bf r}^{*}_{i}-[\,{\bf r}^{*}_{j}+L^{*}\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]}{\big|{\bf r}^{*}_{i}-[\,{\bf r}^{*}_{j}+L^{*}\sum_{l=1}^{3}n_{l}{\bf e}_{l}\,]\big|^{3}}\Bigg\}\>.$
As was already explained above, the sums over $k$ and $n_{k}$ are actually calculated over the cubic shells of increasing radii, composed of the “mirror” cells, until the required accuracy of convergence is achieved.
II.3 The method of integration
@@
Next, they are integrated numerically by the second-order Runge–Kutta method with a constant stepsize. In fact, this method is rather inefficient, and a much better computational performance could be achieved by using the higher-order integrators with adaptive stepsize control. This was done, for example, in our earlier works Dumin and Lukashenko (2022, 2024), where we employed subroutines from the Numerical Recipes library Press et al. (1992). Unfortunately, the adaptive stepsize control operates reliably only for the reflective boundaries and encounters serious problems in the case of periodic boundary conditions (for more details, see Appendix A in paper Dumin and Lukashenko (2022)). So, since the geometrical setup of our simulations (Fig. 1) inherently incorporates the periodic boundaries, we had to use a rather primitive integration scheme.
@@
Yet another very subtle point of our modeling is usage of the exact (singular) Coulomb potentials of the charged particles, without any softening or cut-offs at the small distances, which was a common practice in many earlier papers Lankin and Norman (2009b); Niffenegger, Gilmore, and Robicheaux (2011); Tiwari, Shaffer, and Baalrud (2017). We preferred to avoid such a simplification, because formation of the bound states (resulting in the recombination) is rather sensitive to the specific form of the potential. Particularly, the closed trajectories are possible only in the exact Coulomb field Landau and Lifshitz (1976). (Yet another option for the formation of closed orbits might be the potential of harmonic oscillator, proportional to $r^{2}$; but this is beyond our physical context.)
@@
Unfortunately, if the simulated plasma was initially cold enough, then the electrons begin to fall onto the nearest ions, sometimes resulting in the almost head-on collisions, which cannot be resolved by our integration algorithm. (Such cases can be identified by the sharp jumps in total kinetic energy of the electrons and/or asymmetric profiles of the kinetic energy with respect to a pericenter of the electron orbit.) Then, the integration was canceled, and the corresponding initial conditions were discarded. So, we employed only the “favorable” initial conditions, which do not lead to failures in the very beginning of simulations. In fact, the number of failures was an order of magnitude greater than the number of favorable cases. However, this did not increase dramatically the computational cost of our simulations, because the failures typically occurred at the very early stage of the integration (about a few $\tau$ from the beginning), while we were interested in the long-term dynamics (a few hundred or thousand $\tau$). So, a dozen of failures did not take too much time.
II.4 Initial conditions

### v2

# Non-equilibrium classical recombination in the expanding ultracold plasmas

@@ Abstract
The efficiency of recombination is of crucial importance for the existence of ultracold plasmas (UCP), particularly, the ones formed in the magneto–optical traps. Unfortunately, the equilibrium thermodynamic treatment of the ionization–recombination processes is inappropriate for the evolving UCP clouds, while the straightforward kinetic simulation encounters the problem of huge difference in the spatial and temporal scales for free and bound motion of the electrons. As a result, only the “virtual” electron–ion pairs are usually reproduced in such modeling, and it is necessary to employ some heuristic criteria to identify them with the recombined atoms. It is the aim of this paper to present the first successful ab initio simulation of the non-equilibrium recombination in the evolving UCP plasmas. We employed a special algorithm, which is based on using the “scalable” reference frame, co-moving with the expanding substance. Then, the recombination events are identified by a series of sharp equidistant peaks in the kinetic and/or potential energies, which are caused by the captured electrons passing near the pericenters of their orbits; and this is confirmed by a detailed inspection of their trajectories. Thereby, we were able to trace the real—rather than “virtual”—electron–ion pairs, and the total efficiency of their formation was found to be about 20%, which is in agreement with the laboratory measurements.
@@ I Introduction
The ultracold plasmas (UCP), formed after release of the bunches of photoionized atoms from the magneto-optical traps Killian et al. (1999), represent a relatively new branch of plasma physics, emerged in the very late 1990s and early 2000s Gould and Eyler (2001); Bergeson and Killian (2003); Killian (2007); Killian et al. (2007). The possibility of existence of such plasmas for an appreciable time interval crucially depends on the efficiency of recombination, since it might be intuitively expected that the rarefied charged particles with a sufficiently strong Coulomb coupling will quickly collapse into the neutral pairs.
@@
It is interesting to mention that the first molecular-dynamic simulations of such plasmas started already in the early 1990s, i.e., a decade before the first experiments with UCP. Surprisingly, these simulations demonstrated that the ultracold plasmas experienced a very small recombination, i.e., could survive for a sufficiently long time Mayorov, Tkachev, and Yakovlenko (1994). Unfortunately, it was suspected by some opponents that the anomalous resistance of charged particles with respect to the recombination might be just an artifact of the particular numerical methods. For example, the recombined neural pairs could be efficiently destroyed by the collisions with walls of the computational box Ignatov et al. (1995); Mayorov, Tkachev, and Yakovlenko (1995), if the reflective boundary conditions were imposed there. Moreover, the anomalously small recombination rate might be caused just by the accumulation of numerical errors, preventing formation of the bound electron–ion pairs. This problem is especially severe because of the huge difference in the characteristic temporal and spatial scales for the free and bound motion of the charged particles in plasmas.
@@
In view of the above-mentioned problem, the most of works on recombination in UCP employed the artificial two-step procedure: Firstly, the electron distribution function was simulated numerically by the molecular-dynamic approach, and then it was used for analytical estimates of the recombination rate, e.g., employing the idea of electron diffusion in the energetic space Bobrov et al. (2011).
@@
On the other hand, as regards ab initio modeling of recombination, the most of works employed some additional criteria for the formation of the bound electron–ion pairs: For example, it was assumed that the act of recombination took place if an electron performed a certain number of revolutions about the nearest ion (e.g., 4 complete revolutions) Lankin and Norman (2009a). Yet another approach was to assume that all electrons located closer than a specified distance from the nearest ions (e.g., 20% of the Wigner–Seitz radius) will ultimately recombine Niffenegger, Gilmore, and Robicheaux (2011). These artificial criteria were necessary because the commonly-used numerical schemes did not allow ones to get sufficiently stable electron–ion pairs: they were destroyed rather quickly due to accumulation of the computational errors. Besides, these works usually performed the simulations within a box of a fixed size, which is poorly relevant to the actual experimental conditions, since the UCP clouds typically expand by one or two orders of magnitude in the course of experiments.
@@
It is the aim of the present work to employ a specially designed numerical scheme that is able, firstly, to take into account the considerable degree of expansion of the plasma cloud and, secondly, to trace formation of the stable electron–ion pairs (which should be subsequently converted to neutral atoms due to the quantum processes). In fact, the basic idea of this algorithm was already used in our earlier paper Dumin (2011), which was aimed at ab initio simulation of the electron temperature evolution in UCP bunches released from the magneto-optical traps. The corresponding functional dependence obtained in that work, $T_{e}\propto t^{\alpha}$ with exponent $\alpha\in[-1.25,-1.08]$, was in perfect agreement with the experiment Fletcher, Zhang, and Rolston (2007), where $\alpha=-(1.2{\pm}0.1)$. It is important to emphasize that this coincidence was achieved exactly from “the first principles”, without postulation of any additional heat sources or sinks. As will be shown below, the same numerical scheme provides very good results also in the simulation of recombination.
@@ II.1 Geometrical configuration
Figure 1: Sketch of the plasma cloud (gray segment with a blue boundary) composed of the identical cells of size $L$, uniformly expanding with velocity $u_{0}$.
Following the design of typical UCP experiments, we shall consider a large volume of plasma expanding uniformly in all directions, starting from the center $r=0$. For the sake of simplicity, we assume that the cloud is uniform and has a sharp boundary at $r=R(t)$; see Fig. 1. Next, let this cloud be composed of the identical (or “mirror”) cubic cells filled with some number of ions $N$ and the same number of electrons, and $N_{\rm m}$ be the number of cells from the center to the boundary. Each cell expands with velocity $u_{0}$ with respect to its opposite face. So, at the cloud boundary $R=N_{\rm m}L$, the velocity becomes $V=N_{\rm m}\,u_{0}$. In fact, this scheme closely resembles the idea of Hubble-Lemaître expansion, commonly used in cosmology Zel’dovich and Novikov (1983).
It is well known that the initial thermal energy of plasma (which, in the case of photoionization, is concentrated mostly in electrons) quickly transforms into kinetic energy of macroscopic motion of the cloud; e.g., papers Vikhrov et al. (2020, 2021) and references therein. In other words, $m_{i}\,V^{2}\sim\,k_{\rm B}\,T_{e0}$, where $T_{e0}$ is the initial electron temperature, $V$ is the velocity of the cloud boundary, $m_{i}$ is the ionic mass, and $k_{\rm B}$ is Boltzmann constant. Consequently,
$V\sim\sqrt{k_{\rm B}\,T_{e0}/m_{i}}$ (1)
(we did not write here any numerical coefficient, since it depends substantially on the distribution of density inside a more realistic cloud). Just this formula can be used to link the experimental quantities with the simulation parameters.
II.2 Equations of motion
Figure 2: Sketch of calculation of the Coulomb sums, where each electron $e$ interacts not only with a specified ion $i$ inside the basic cell (solid arrow) but also with infinite number of its “mirror” images $i^{\prime}$ (dashed arrows).
The simulations are based on solving the classical equations of motion for $N$ electrons inside the “basic” cell:
$m_{e}\,\frac{d^{2}{\rm\bf r}_{i}}{dt^{2}}=\,{\rm\bf F}_{i}\,$, (2)
where ${\rm\bf r}_{i}$ are the electron coordinates, and $m_{e}$ is the electron mass. The Coulomb force ${\rm\bf F}_{i}$, acting on the $i$’th electron, is calculated taking into account all particles in the system i.e., not only within the same basic cell but also the all its “mirror” images (Fig. 2):
@@ II.2 Equations of motion
$+$ $\sum_{{\bf n}\,\in\,\mathbb{Z}^{3}\backslash\{{\bf 0}\}}\Bigg\{\sum_{j=1}^{N}e^{2}\frac{({\bf R}_{j}+L\,{\bf n})-{\bf r}_{i})}{\big|({\bf R}_{j}+L\,{\bf n})-{\bf r}_{i}\big|^{3}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\>+\;\sum_{j=1}^{N}\,e^{2}\frac{{\bf r}_{i}-({\bf r}_{j}+L\,{\bf n})}{\big|{\bf r}_{i}-({\bf r}_{j}+L\,{\bf n})\big|^{3}}\Bigg\}\>.$
[…]
@@
In fact, summation in formula (3) is performed over a finite number of mirror cells until the specified accuracy of convergence of the Coulomb forces is achieved. These mirror cells are arranged into the cubic shells of increasing radii around the basic cell, as illustrated in Fig. 3. Strictly speaking, the sums like (3) are conditionally convergent, i.e., the result can depend on the summation order. The method of summation depicted in the figure seems most reasonable from the physical point of view. Really, effect of the mirror cells located at the larger radii should be obscured by some additional physical perturbations, thereby providing a “physical regularization” of the conditionally-convergent Coulomb sums.
Although we integrated the equations of motion for a relatively small number of charged particles in the basic cell (e.g., a few dozens or a few hundreds), the total number of particles taken into account in calculation of the Coulomb sums was huge. For example, if the relative accuracy of calculations was specified to be about $10^{-4}$ or $10^{-5}$, then even at $N=10$ the total number of particles involved in the computation turned out to be from a few hundred thousand to a few million. Let us emphasize that this number was not fixed but chosen “dynamically” at each step of integration.
In fact, the same type of summation as in formula (3) is commonly implemented in the condensed-matter physics by the so-called Ewald method. (Its original idea is outlined, e.g., in textbook Ziman (1972), and a more elaborated version can be found in paper Demyanov and Levashov (2022).) Most probably, the same approach can be applied also to the problem under consideration, but this issue requires a further study.
II.3 The scalable reference frame
Yet another serious problem in simulation of the expanding plasma clouds is a considerable change in the spatial scale of the system—and, therefore, in the Coulomb forces—in the course of time, which makes it difficult to keep the accuracy of integration at a stable level. To get around this obstacle, we introduce the scalable reference frame, expanding with the average expansion rate of the plasma. Namely, length of the basic cell is assumed to increase as
$L(t)=L_{0}+u_{0}\,t,$ (4)
@@ II.3 The scalable reference frame
In other words, the simulated particles perform their motion in a box of the fixed size (9), thereby mitigating the problem of large variation of the spatial coordinates in the case of considerable plasma expansion.
As usual, the periodic boundary conditions are imposed at the box boundaries: when one of the particles leaves the box through one of its faces, the identical particle with the same velocity enters this box through the opposite face.
The equation of motion of the $i$’th electron (2) in the scale-dependent dimensionless variables is reduced to
@@
$+$ $\sum_{{\bf n}\,\in\,\mathbb{Z}^{3}\backslash\{{\bf 0}\}}\Bigg\{\sum_{j=1}^{N}\frac{({\bf R}^{*}_{j}+L^{*}\,{\bf n})-{\bf r}^{*}_{i})}{\big|({\bf R}^{*}_{j}+L^{*}\,{\bf n})-{\bf r}^{*}_{i}\big|^{3}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\>+\;\sum_{j=1}^{N}\frac{{\bf r}^{*}_{i}-({\bf r}^{*}_{j}+L^{*}\,{\bf n})}{\big|{\bf r}^{*}_{i}-({\bf r}^{*}_{j}+L^{*}\,{\bf n})\big|^{3}}\Bigg\}\>.$
As was already explained above, the sums over $\bf n$ are actually calculated over the cubic shells of increasing radii, composed of the “mirror” cells, until the required accuracy of convergence is achieved.
II.4 The method of integration
@@
Next, they are integrated numerically by the second-order Runge–Kutta method with a constant stepsize. In fact, this method is rather inefficient, and a much better computational performance could be achieved by using the higher-order integrators with the adaptive stepsize control. This was done, for example, in our previous works Dumin and Lukashenko (2022, 2024), where we employed the subroutines from Numerical Recipes library Press et al. (1992). Unfortunately, the adaptive stepsize control operates reliably only for the reflective boundaries and encounters serious problems in the case of periodic boundary conditions (for more details, see Appendix A in paper Dumin and Lukashenko (2022)). So, since the geometrical setup of our simulations (Fig. 2) inherently incorporates the periodic boundaries, we were enforced to use here a rather primitive integration scheme.
@@
Yet another very subtle point in our modeling is usage of the exact (singular) Coulomb potentials of the charged particles, without any “softening” or cut-offs at the small distances, which was a common practice in many earlier papers Lankin and Norman (2009b); Niffenegger, Gilmore, and Robicheaux (2011); Tiwari, Shaffer, and Baalrud (2017). We preferred to avoid such a simplification, because formation of the bound states (resulting in the recombination) is rather sensitive to the specific form of the potential. Particularly, the closed trajectories are possible only in the exact Coulomb field Landau and Lifshitz (1976). (Yet another option for the formation of closed orbits is the potential of harmonic oscillator, proportional to $r^{2}$; but this is evidently beyond our physical context.)
@@
Unfortunately, if the simulated plasma was initially cold enough, then the electrons began to fall onto the nearest ions, sometimes resulting in the almost head-on collisions, which could not be resolved by our integration algorithm. Such cases were identified by the sharp jumps in the kinetic energy of electrons and their asymmetric temporal profiles (with respect to the passage of pericenters); for the particular examples, see Appendix A in the present paper. Then, the integration was cancelled, and the corresponding initial conditions were discarded. So, we used only the “favorable” initial conditions, which did not lead to the failures. In fact, the number of failures was a few times greater than the number of successful attempts. Nevertheless, this did not increase appreciably the computational cost of our simulations, because the failures typically occurred at the very early stage of integration (about a few $\tau$ from the beginning), while we were interested in the long-term dynamics (several tens or hundreds $\tau$). So, a few failures per one successful attempt did not take too much time.
II.5 Initial conditions